Expand and simplify (2-(1/n))^6
Thanks

Answer by fractalier(6550)

We can use the tenets of binomial expansion to help us...so that
(2-(1/n))^6 = 2^6 - 6*(2^5)*(1/n) + 15*(2^4)*(1/n)^2 - 20*(2^3)*(1/n)^3 + 15*(2^2)*(1/n)^4 - 6*(2)*(1/n)^5 + (1/n)^6 =
64 - 192/n + 240/n^2 - 160/n^3 + 60/n^4 - 12/n^5 + 1/n^6